find the general solution of the differential equation x dy/dx + 4y =x^3


1
Expert's answer
2021-04-29T11:50:34-0400

Let us find the general solution of the differential equation "x \frac{dy}{dx} + 4y =x^3". For this let us multiply both parts by "x^3:" "x^4y' + 4x^3y =x^6". The latter equation is equivalent to "(x^4y)'=x^6". It follows that "x^4y=\frac{x^7}{7}+C," and hence the general solution is "y=\frac{x^3}{7}+\frac{C}{x^4}."